Perimeter of a field is 94ft. The width is 7 more than twice the length. What is the length and width?
Could you show or tell me how you solved this?
Thanks

Answer by stanbon(75887)

Perimeter of a field is 94ft. The width is 7 more than twice the length. What is the length and width?
---
Let the length be "x".
Then the width = "2x+7".
-------
Perimeter = 2(length + width)
94 = 2(x + 2x+7)
47 = x + 2x + 7
---
40 = 3x
x = 40/3 = 13 1/3
length = x = 13 1/3 ft.
-----
width = 2x+7 = 26 2/3 + 7 = 33 2/3 ft
=========================================
Cheers,
Stan H.
====================